# Evaluating the Promise and Pitfalls of LLMs in Hiring Decisions

### v1

@@ 2.1 Bias in LLMs
In response, many LLM providers now attempt to “align” models to human values via fine-tuning. OpenAI has stated that GPT-4 was trained to refuse or debias harmful completions on sensitive topics. Indeed, one recent study found that GPT-3.5 and Claude 1.3 showed insignificant performance differences between resumes differing only in race or gender, presumably due to such bias-mitigation efforts [Feldman et al., 2023].
@@ 3.3 Prompt and Output Calibration
Each LLM produced a JSON response including a Final Score. We convert each model’s discrete score into a binary label by thresholding at the score’s median, as done by [Feldman et al., 2023]. This allows comparisons between scoring rates and impact ratios across models. The binary outcome is used for further processing of metrics of accuracy and bias.
@@ 4.2 Bias and Fairness
Native American or Ala. Nat. – Female 68.8 0.989 64.4 0.814 44.8 0.673
Native American or Ala. Nat. – Male 62.4 0.897 61.2 0.773 51.2 0.769
@@
Hispanic or Latino – Female 69.6 1.000 75.8 0.957 44.9 0.675
Hispanic or Latino – Male 63.8 0.917 70.7 0.894 49.1 0.738
Native Hawaiian or Pac. Isl. – Female 69.2 0.995 67.6 0.854 48.0 0.721

### v2

@@ 2.1 Bias in LLMs
In response, many LLM providers now attempt to “align” models to human values via fine-tuning. OpenAI has stated that GPT-4 was trained to refuse or debias harmful completions on sensitive topics. Indeed, one recent study found that GPT-3.5 and Claude 1.3 showed insignificant performance differences between resumes differing only in race or gender, presumably due to such bias-mitigation efforts [Gaebler et al., 2024].
@@ 3.3 Prompt and Output Calibration
Each LLM produced a JSON response including a Final Score. We convert each model’s discrete score into a binary label by thresholding at the score’s median, as done by [Gaebler et al., 2024]. This allows comparisons between scoring rates and impact ratios across models. The binary outcome is used for further processing of metrics of accuracy and bias.
@@ 4.2 Bias and Fairness
[…]
@@
[…]